OpenStudy (anonymous):

So I have to solve for |4w+9|=|4w+5| I concluded that 0=4 therefore there is no solution, am I right?

OpenStudy (anonymous):

let see you have four possible equation here

OpenStudy (anonymous):

(4w+9)=(4w+5) (4w+9)=-(4w+5) - (4w+9)=(4w+5) - (4w+9)=-(4w+5)

OpenStudy (anonymous):

So i would solve for each equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

0=4 would be one of them?

OpenStudy (anonymous):

w=-7/4 would be the second?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the first solution would be no solution right?

OpenStudy (anonymous):

yes , first and fourth